OpenStudy (anonymous):

if the radius of a circle is increased by 4 inches, the area of the resulting circle is 100pi in^2. Fin the radius of the original cricle,

OpenStudy (anonymous):

A = pi * r^2 We can let r = x + 4 since the radius has been increased by 4. Where does the 100 pi go?

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

Hmmmm ..... how about area?

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

100 pi = pi * (x+4)^2 then divide both sides by pi 100 = (x+4)^2 All is good?